OpenStudy (anonymous):

The local amusement park charges $16.95 per person. However, after 5 PM the charge is reduced to $12.55 per person. A group of 8 adults is planning to go to the park. How much would the save if they go after 5 PM???

OpenStudy (anonymous):

@jut21 could you help out?

OpenStudy (lynfran):

well if they go after 5pm and since its 8 of them then 8*12.55=?

OpenStudy (anonymous):

100.40???

OpenStudy (lynfran):

and since the local charges was 16.95 we subtract that from 100.40 we get?

OpenStudy (lynfran):

or another way to look at it..16.95*8=135.6 and 12.55*8=100.4 now 135.6-100.4=

OpenStudy (lynfran):

?

OpenStudy (lynfran):

ok they would save $35.2